# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

### v3

@@ 3.4. RQ2: Effects of Unseen Class Ratio
To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories underwent a progressive reduction during the training process. The results are visually depicted in Fig.  3. Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%-¿40% and 20%-¿60% respectively. We conclude the following conclusions based on the outcome results,
  • All deep hashing methods manifest a discernible decline in performance concomitant with an escalation in the number of unseen classes on three datasets. Notably, this performance deterioration not only intensifies but also amplifies in magnitude, as evidenced by both the MAP@all and AUC metrics.
  • COMAE exhibits a performance superiority across various unseen class ratios, along with the most minimal decrease.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\epsilon$
Figure 6. The effects of hyper-parameters on AWA2 dataset.
(a) COMAE
(b) HashNet
(c) DPN
(d) OrthoCos
(e) CSQ
(f) BiHalf
Figure 7. t-SNE visualizations of hashing code on test unseen classes. The 10 colors denote 10 different unseen classes.
3.5. RQ3: Distance Visualization of Intra and Inter Classes
@@
To verify whether the classwise constraint can broaden the distance of inter-class and reduce the distance of intra-class, we provide a quantitative comparison of the distance between the mean value of intra-class distances and inter-class distances, under the setting of without loss $\mathcal{L}_{\text{contrastive}}$ and with the loss. Fig. 5 summarizes the histogram of intra-class and inter-class distances on the AWA2 dataset. We adjust the frequency normalization so the total of all histogram bins equals one. The directional marker indicates the separability in Hamming distances, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$. Despite similar distributions for inter-class distances, the distinction between the average intra-class distance (indicated by the purple dotted line) and the average inter-class distance (shown by the orange dotted line) in scenario (b) exceeds that in scenario (a). This approach visually emphasizes the effect of the classwise constraint on class separability metrics. Quantificationally, the difference in (a) is 14.38, and the difference in (b) is 15.50. The results show that our proposed classwise constraint can enlarge the distance of the inter-class and reduce the distance of the intra-class.
3.6. RQ3: Component Analysis
@@
To provide further insight into COMAE, we conduct component analysis on AWA2 and CUB datasets with 64-bit hashing-code lengths. We configure several variants of COMAE as a) COMAE-V1 is the COMAE model without $\mathcal{L}_{\text{pointwise}}$; b) COMAE-V2 represents the COMAE without $\mathcal{L}_{\text{pairwise}}$; c) COMAE-V3 refers to the COMAE without $\mathcal{L}_{\text{classwise}}$. Results are shown in Table 2. Compared with other components, the pointwise objective plays the most significant role in AWA2, and the pairwise objective in CUB. Also, the performance of COMAE on both datasets surpasses that of variants, validating the necessity of each component.
@@
Table 2. The mAP@all and AUC results with switching model components in different datasets.
AWA2 CUB
mAP@all AUC mAP@all AUC
COMAE-V1 0.4830 0.4781 0.2069 0.2418
COMAE-V2 0.4857 0.4791 0.1867 0.2240
COMAE-V3 0.4835 0.4782 0.2064 0.2412
COMAE 0.4883 0.4812 0.2094 0.2416
3.7. RQ4: Parameter Analysis
We analyze the impact of critical hyper-parameters, i.e., $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ that influence the weight of the auxiliary losses in Eqn. 9 and the threshold $\epsilon$ for selecting positive samples in Eqn. 4. These evaluations were performed using the AWA2 dataset with 64-bit hash codes, and the results are detailed in Fig.  6. As the increase of hyper-parameters, the performance of our model initially rises and then decreases. Hence, we choose the best hyper-parameters corresponding to the point at peak performance.
@@ 4. Conclusion
In this paper, we present COMAE for a comprehensive attribute exploration in ZSH. COMAE depicts the relationships between seeable and unseeable classes through meticulously designed explorations: point-wise, pair-wise, and class-wise constraints. Our model employs an attribute prototype network to regress attributes, facilitating the learning of local features. Then, we employ contrastive learning to capture the context of attributes, moving beyond instance-independent optimization. Finally, we introduce the class-wise constraint to cohesively enhance hashing learning, image representations, and visual attributes. Experiments demonstrate the excellent performance of COMAE.

### v5

@@ 4.3. RQ2: Effects of Unseen Class Ratio
To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories undergoes a progressive reduction during the training process. The results are visually depicted in Fig.  3. Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%-¿40% and 20%-¿60% respectively. Based on the outcome results, we conclude the following conclusions: (1) All deep hashing methods show a noticeable decline in performance along with an increase in the number of unseen classes on three datasets. Notably, this performance drop not only intensifies but also amplifies in magnitude, as evidenced by both the MAP@all and AUC metrics. (2) COMAE exhibits a performance superiority across various unseen class ratios, with minimal decrease.
Figure 5. Case study on the CUB dataset.
Table 2. The average running time for each epoch
Method AWA2 CUB SUN
mAP Train Infer. mAP Train Infer. mAP Train Infer.
PIXEL 0.5456 483s 0.154s 0.2819 625s 0.183s 0.1741 568s 0.231s
COMAE 0.5801 45s 0.011s 0.2532 48s 0.052s 0.1764 47s 0.075s
4.4. RQ3: Case Study
We illustrate the performance of COMAE through case studies with 64-bit hash codes, as shown in Fig 5. We present the top 10 retrieved images for each query using COMAE (first row) and PIXEL (second row). Correctly retrieved images are outlined in green, while incorrect ones are outlined in red. COMAE demonstrates superior retrieval performance, effectively capturing key visual features. For queries with distinctive visual characteristics, such as the yellow-headed blackbird or the blue bunting, COMAE retrieves highly accurate results, primarily by leveraging attributes like feather color, body shape, and posture. In more challenging cases involving birds with subtle differences or overlapping features, such as cuckoos and flycatchers, COMAE still maintains strong retrieval accuracy but occasionally retrieves images with semantically similar attributes. For instance, brownish birds with similar body structures sometimes lead to false positives, as observed in some incorrect retrievals.
Overall, for intact retrieval targets, COMAE achieves an accuracy of 80%, while in complex cases involving partial occlusion, highly similar species, or ambiguous visual features, COMAE retains a 70% accuracy rate, outperforming PIXEL in zero-shot hashing tasks. The comparison between COMAE and PIXEL highlights COMAE’s robustness in learning effective feature representations and reducing semantic drift in image retrieval.
4.5. RQ4: Efficiency Analysis
We conduct model efficiency analysis with a 64-bit hashing length in this section. Table 2 reports the average training time per epoch and inference time per epoch. We select the best and latest baseline, PIXEL, as our comparable baseline. We find that COMAE achieves competitive computational efficiency. In the largest dataset, SUN, COMAE reduces the training time by over 91.7%, and the inference time is reduced from 0.231s (PIXEL) to 0.075s (COMAE), with an improvement of 67.5%. This demonstrates that COMAE not only trains significantly faster but also achieves efficient inference, making it well-suited for real world large-scale retrieval systems.
4.6. RQ5: Model Analysis
@@
Distance Visualization of Intra and Inter Classes. For model analysis, we first conduct the distance visualization of intra and inter classes, to verify whether the classwise constraint can broaden the distance of inter-class and reduce the distance of intra-class. We provide a quantitative comparison of the distance between the mean value of intra-class distances and inter-class distances, under the setting of without the contrastive loss $\mathcal{L}_{\text{pairwise}}$ and with it. Fig. 4 summarizes the histogram of intra-class and inter-class distances on the AWA2 dataset. We adjust the frequency normalization so that the total of all histogram bins equals one. The directional marker indicates the separability in Hamming distances, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$. Despite similar distributions for inter-class distances, the distinction between the average intra-class distance (indicated by the purple dotted line) and the average inter-class distance (shown by the orange dotted line) in scenario (b) exceeds that in scenario (a). This approach visually emphasizes the effect of the classwise constraint on class separability metrics. Quantificationally, the difference in (a) is 14.38, and the difference in (b) is 15.50. The results show that our proposed classwise constraint can enlarge the distance of the inter-class and reduce the distance of the intra-class.
@@
Component Analysis. To provide further insight into COMAE, we conduct component analysis on AWA2 and CUB datasets with 64-bit hashing-code lengths. We configure several variants of COMAE as (1) COMAE-V1 is the COMAE model without $\mathcal{L}_{\text{pointwise}}$; (2) COMAE-V2 represents the COMAE without $\mathcal{L}_{\text{pairwise}}$; (3) COMAE-V3 refers to the COMAE without $\mathcal{L}_{\text{classwise}}$. Results are shown in Table 3. Compared with other components, the pointwise objective plays the most significant role in AWA2, and the pairwise objective in CUB. Also, the performance of COMAE on both datasets surpasses that of the variants, validating the necessity of each component.
@@
Table 3. The mAP@5000 and AUC results with switching model components in different datasets.
AWA2 CUB SUN
mAP@5000 AUC mAP@5000 AUC mAP@5000 AUC
COMAE-V1 0.5093 0.4364 0.1988 0.2333 0.1420 0.1286
COMAE-V2 0.5122 0.4981 0.1794 0.2161 0.1310 0.1243
COMAE-V3 0.5099 0.5025 0.1983 0.2327 0.1350 0.1264
COMAE 0.5149 0.5106 0.2012 0.2331 0.1465 0.1302
@@ 5. Conclusion
Zero-Shot Hashing stands out for its high efficiency and robust generalization, particularly in large-scale retrieval scenarios. In this paper, we present COMAE for a comprehensive attribute exploration in ZSH. COMAE depicts the relationships between seeable and unseeable classes through meticulously designed explorations: point-wise, pair-wise, and class-wise constraints. Furthermore, theoretical analysis is provided to show the effectiveness of COMAE. Our model employs an attribute prototype network to regress attributes, facilitating the learning of local features. Then, we employ contrastive learning to capture the context of attributes, moving beyond instance-independent optimization. Finally, we introduce the class-wise constraint to cohesively enhance hashing learning, image representations, and visual attributes. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes.
Acknowledgements.
We would like to thank Zeyu Dong, Ning Cao, Shuai Liu, Pengfei Wang, and Yuchen Yan who discussed with us and provided instructive suggestions. This work is partially supported by the National Natural Science Foundation of China (No.92470204), the Beijing Natural Science Foundation (No.4254089), the Postdoctoral Fellowship Program of CPSF (No.GZC20232736), and the China Postdoctoral Science Foundation Funded Project (No.2023M743565).